# Tree-Guided Diffusion Planner

### v1

@@ 5.1 Baselines and Ablations
  • Diffuser Janner et al. (2022):   Diffusion-based approach that plans by iteratively denoising complete trajectories.
  • AdaptDiffuser Liang et al. (2023):   Fine-tune pretrained Diffuser with synthetic expert demonstrations.
@@
  • Monte-Carlo Sampling with Selection (MCSS) Lu et al. (2025):   Sample multiple trajectories from the pretrained Diffuser and select the best trajectory based on the guidance score. Monte Carlo sampling methods effectively explore the solution space and approximate optimal trajectories Shapiro (2003).
  • Stochastic Sampling (MCSS+$S$) Wang et al. (2025):   MCMC-based, training-free guided planning method requiring $M$ times more computation than MCSS, where $M$ is the number of iterations in the inner loop of diffusion sampling. Following Wang et al. (2025), we set $M=4$ in all our experiments.
@@ 5.3 KUKA Robot Arm Manipulation
We evaluate the test-time planning performance of TDP and baselines on robotic arm manipulation tasks. Diffusion planners are pretrained on arbitrary block stacking demonstrations collected from PDDLStream Garrett et al. (2020), and are typically evaluated on downstream tasks such as conditional stacking Janner et al. (2022) and pick-and-place Liang et al. (2023), where test-time goals are specified to the planner. Both tasks aim to place randomly initialized blocks into their corresponding target locations in a predetermined order. Pick-and-place task is more challenging, as it involves placing each block at a unique target without access to expert demonstrations. To better isolate test-time planning capability, we evaluate a variant of the conditional stacking task Liang et al. (2023) without the use of the pretrained classifier with expert data. We refer to this variant as PnP (stack), and the original pick-and-place task is denoted as PnP (place). For both tasks, the pretrained diffusion planner is guided to generate 4 sequential manipulation trajectories — one for each block — toward its target location $\boldsymbol{s}_{\text{target}}$, using a naïve guidance objective defined as $\mathcal{J}(\boldsymbol{\tau})=-\sum_{i}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{target}}\right\rVert$. As shown in Table˜2, TDP achieves an average improvement of 10% over MCSS and 20% over TAT in two PnP tasks, demonstrating strong generalization to diverse task configurations. Notably, TDP (w/o PG) outperforms MCSS by 18% on PnP (place), underscoring the role of our sub-trajectory refinement mechanism in such out-of-distribution planning scenarios.
@@
Moreover, we carefully design a more challenging test-time manipulation task which extends PnP (place), namely pick-and-where-to-place (PnWP). As shown in Fig. 4, the agent must find a global optimal trajectory given non-concave guide function $\mathcal{J}(\boldsymbol{\tau})=-\sum_{i}c_{1}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{local}}\right\rVert-c_{2}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{mid}}\right\rVert+c_{3}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{global}}\right\rVert$ where $c_{1},c_{2},c_{3}$ are positive constants and $\boldsymbol{s}_{\text{local}},\boldsymbol{s}_{\text{mid}},\boldsymbol{s}_{\text{global}}$ are the local, middle, and global target positions, respectively. Both $\boldsymbol{s}_{\text{global}}$ and $\boldsymbol{s}_{\text{local}}$ are equidistant from the robot arm’s attachment point. Since the local optimum has a wide peak while the global optimum has a narrow peak, agents easily get trapped in local optima without sufficient exploration. While PnP requires fitting blocks into a target configuration, PnWP challenges planners to distinguish globally optimal arrangements from suboptimal ones.
@@
We report the performance of TDP, baselines and ablations in Table˜2. Mono-level guided sampling methods (i.e., AdaptDiffuser, MCSS(+SS), TAT, and TDP (w/o child)) tend to converge to local optima, often stacking all blocks at a single position, since a landscape of local optima is spread out in a broader range as shown in Fig. 4. In contrast, bi-level sampling approaches (i.e., TDP and TDP (w/o PG)), which combine parent branching and sub-trajectory refinement, are better able to consistently identify globally optimal placements. Notably, TDP outperforms AdaptDiffuser both on the standard benchmark (PnP) and on the custom task (PnWP). This demonstrates that TDP’s test-time scalability and generalization enable zero-shot planning that surpasses the training-per-task approach.
@@
The key advantage of our method over the baselines comes from employing unconditional PG in the parent branching phase, enabling broad exploration of the trajectory space. Both TDP and its ablations produce diverse trajectories (see Fig. 5-a); however, TDP (w/o child) often fails to find the global optimal placement due to the use of conditional PG in parent branching, which can bias samples toward local optima. The mean pairwise trajectory distance decreases when PG is combined with sub-tree expansion, as the generated sub-trajectories share segments with their corresponding parent trajectories. Unconditional PG enables generating diverse, gradient-free parent trajectories. When combined with sub-tree expansion guided by the objective, this bi-level strategy supports compositional solutions that successfully stack blocks into global optimal placement (see Fig. 5-b). More detailed information on the evaluation metrics among these tasks is available in Appendix˜F.
@@ 5.4 AntMaze Multi-goal Exploration
We finally evaluate test-time multi-goal exploration capability on AntMaze Fu et al. (2021), which is more challenging than Maze2D due to its high-dimensional observation space for controlling the embodied agent. Its complexity causes the pretrained diffusion planner to predict shorter horizons than the full trajectory length Dong et al. (2024). We design a new multi-goal task in which the agent must visit all goal positions in the correct order of priority, as specified by the guide function $\mathcal{J}(\boldsymbol{\tau})=\sum_{g\in\mathcal{G}}h_{g}\cdot\exp{(-\sum_{i\leq T_{\text{pred}}}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{g}\right\rVert^{2}/\sigma_{g}^{2})}$ where $\mathcal{G}$ is the set of goal positions and $h_{g}$, $\sigma_{g}$ are parameters for gaussian guide function at each goal $g\in\mathcal{G}$. For example, the agent in Fig. 6 first visits goal $g_{2}$ at $t=t_{3}$. If it subsequently visits $g_{1}$, $g_{4}$, and $g_{3}$ after $t=t_{3}$, it successfully reaches all four goals in the sequence $g_{2}\rightarrow g_{1}\rightarrow g_{4}\rightarrow g_{3}$. However, two precedence rules are violated ($g_{2}\rightarrow g_{1}$, $g_{4}\rightarrow g_{3}$) while the remaining four ($g_{2}\rightarrow g_{4}$, $g_{2}\rightarrow g_{3}$, $g_{1}\rightarrow g_{4}$, and $g_{1}\rightarrow g_{3}$) are satisfied. In this scenario, the agent achieves a goal completion score of 4/4 but only 4/6 priority sequence match accuracy. Maximum accuracy of 6/6 can only be achieved by visiting all goals in the correct prioritized order.
@@ 6 Conclusion
In summary, we propose TDP, a flexible test-time planning framework that leverages a pretrained diffusion planner through a bi-level trajectory sampling process without training. By balancing trajectory diversity and gradient-guided refinement through a branching structure of sampled trajectories, our method addresses key limitations of conventional test-time guided planning. Empirical results across both structured and compositional manipulation tasks demonstrate consistent performance gains over existing baselines, particularly in scenarios that demand out-of-distribution generalization. Our experiments also highlight the robustness of the framework in handling non-convex, multi-objective guidance and non-differentiable constraint problems, where naïve gradient guided methods often fail.
@@
While TDP outperforms existing planning approaches in the suite of challenging test-time control tasks, our bi-level trajectory generation process introduces additional computational cost due to the expanded search in trajectory space and the pairwise trajectory distance calculation. We analyze the additional computational time required for the two PnP tasks and the PnWP task in Appendix˜G. Future work may explore more efficient search strategies or learned priors to reduce overhead while ensuring sufficient exploration and preserving planning performance.
@@ Appendix C Proof of Proposition 1
where the conditional distribution $X_{t}|X_{0}$ in forward process is Gaussian, i.e., $\mathcal{N}(\alpha(t)x_{0},h(t)I)$ where $h(t)=1-\alpha^{2}(t)=1-\exp(-\sqrt{t})$ Guo et al. (2024). First we consider the case when $X_{0}\sim\mathcal{N}(0,I)$. The orthogonal gradient can be approximated as $\nabla\mathcal{J}_{\perp}(X)\approx-\frac{1}{\sigma_{2}^{2}}(X_{\perp}-w_{\perp})\mathcal{J}_{2}(X)$ by assumption. As $\mathcal{J}_{2}(X)\geq\mathbb{E}_{X_{0}}[\mathcal{J}_{2}(X_{0})]=:b_{0}>0$ and taking expectation, it is derived by linear ODE $d\mathbb{E}[X_{t,\perp}]=\left[\left(\gamma(t)-\frac{b_{0}}{\sigma_{2}^{2}}\right)\mathbb{E}[X_{t,\perp}]+\frac{b_{0}}{\sigma_{2}^{2}}w_{\perp}\right]dt$. Solving this ODE, we get lower bound of the expectation of displacement of the final state $\mathbb{E}[|X_{T,\perp}|]\geq\exp{\left(-\Phi(T,0)\right)}\left(\frac{b_{0}}{\sigma_{2}^{2}}\int_{0}^{T}\exp{\left(\Phi(s,0)\right)}ds\right)|w_{\perp}|$ where $\Phi(t,s)=\int_{s}^{t}\left(\gamma(u)-\frac{b_{0}}{\sigma_{2}^{2}}\right)du$. Since the coefficient of direction vector $w_{\perp}$ is a tractable positive constant, the final state is always off-subspace. The final state converges linearly as the given guide function satisfies the PL condition and the Lipschitz property Karimi et al. (2020). However, the global optimal solution lies in the subspace spanned by $A$, guided sample converges to local optimal solution which is orthogonal to the subspace.
Second, we consider the case when $X_{0}\sim q_{N_{f}}(\hat{X}^{T},\Sigma^{T})$. As the small perturbation is added to the unconditional sample in the subspace, the orthogonal gradient can be approximated as $\nabla\mathcal{J}_{\perp}(X)\approx-\frac{1}{\sigma_{1}^{2}}(I-AA^{\top})(X-Av_{1})\mathcal{J}_{1}(X)=-\frac{1}{\sigma_{1}^{2}}X_{\perp}\mathcal{J}_{1}(X)$. From Equation˜7, we get linear ODE $d\mathbb{E}[X_{t,\perp}]=\left[\left(\gamma(t)-\frac{1}{\sigma_{1}^{2}}\mathcal{J}_{1}(X)\right)\mathbb{E}[X_{t,\perp}]\right]dt$. As $\mathbb{E}[X_{0,\perp}]=0$, we have $\mathbb{E}[X_{T,\perp}]=0$, which means the final state is on-subspace. Similar to the previous case, the final state converges linearly to the global optimal solution lying on the subspace as the given guide function satisfies the PL condition and Lipschitz property Karimi et al. (2020).
@@ I.2 AntMaze Multi-goal Exploration
TDP surpasses sequential approaches (i.e., MCTD Yoon et al. (2025), Diffusion-Forcing Chen et al. (2024a)) on the standard maze offline benchmarks Park et al. (2025).
@@
TDP has advantages over sequential approaches even in single-task scenarios, in terms of trajectory optimality. For instance, in standard maze benchmarks, the shortest-path trajectory between initial and goal states represents the globally optimal solution, while longer trajectories correspond to suboptimal (local optimal) solutions Chen et al. (2024b). As maze environments scale up, identifying globally optimal solutions requires evaluating combinatorially diverse trajectory candidates with varying structures.
Single-step exploration methods like MCTD optimize actions one step at a time (via guided vs. non-guided action selection), which tends to lead to locally greedy decisions Browne et al. (2012) and results in longer trajectories than the shortest path. In contrast, TDP employs a multi-step exploration framework via bi-level search, enabling per-step action selection based on multi-step future rewards evaluated across diverse-branched trajectory candidates. This approach allows TDP to be able to assess long-term action consequences and avoid the local optimal path.
@@
Learned potentials Corso et al. (2024) offer slight improvements in both sample diversity and overall performance. The results below compare TDP using learned versus fixed potentials on the KUKA benchmarks.

### v2

@@ 5.1 Baselines and Ablations
  • Diffuser [janner2022planningdiffusionflexiblebehavior]:   Diffusion-based approach that plans by iteratively denoising complete trajectories.
  • AdaptDiffuser [liang2023adaptdiffuserdiffusionmodelsadaptive]:   Fine-tune pretrained Diffuser with synthetic expert demonstrations.
@@
  • Monte-Carlo Sampling with Selection (MCSS) [lu2025what]:   Sample multiple trajectories from the pretrained Diffuser and select the best trajectory based on the guidance score. Monte Carlo sampling methods effectively explore the solution space and approximate optimal trajectories [SHAPIRO2003353].
  • Stochastic Sampling (MCSS+SS) [wang2025inferencetimepolicysteeringhuman]:   MCMC-based, training-free guided planning method requiring $M$ times more computation than MCSS, where $M$ is the number of iterations in the inner loop of diffusion sampling. Following [wang2025inferencetimepolicysteeringhuman], we set $M=4$ in all our experiments.
@@ 5.3 KUKA Robot Arm Manipulation
We evaluate the test-time planning performance of TDP and baselines on robotic arm manipulation tasks. Diffusion planners are pretrained on arbitrary block stacking demonstrations collected from PDDLStream [garrett2020pddlstreamintegratingsymbolicplanners] and are typically evaluated on downstream tasks such as conditional stacking [janner2022planningdiffusionflexiblebehavior] and pick-and-place [liang2023adaptdiffuserdiffusionmodelsadaptive], where test-time goals are specified to the planner. Both tasks aim to place randomly initialized blocks into their corresponding target locations in a predetermined order. The pick-and-place task is more challenging because it requires placing each block at a unique target without access to expert demonstrations. To better isolate test-time planning capability, we evaluate a variant of the conditional stacking task [liang2023adaptdiffuserdiffusionmodelsadaptive] without using the pretrained classifier on expert data. We refer to this variant as PnP (stack), and the original task is denoted as PnP (place). For both tasks, the pretrained diffusion planner is guided to generate 4 sequential manipulation trajectories, one for each block, toward its target location $\boldsymbol{s}_{\text{target}}$, using a naïve guidance objective defined as $\mathcal{J}(\boldsymbol{\tau})=-\sum_{i}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{target}}\right\rVert$. As shown in Table˜2, TDP achieves an average improvement of 10% over MCSS and 20% over TAT in two PnP tasks, demonstrating strong generalization to diverse task configurations. Notably, TDP (w/o PG) outperforms MCSS by 18% on PnP (place), underscoring the role of our sub-trajectory refinement mechanism in such out-of-distribution planning scenarios.
@@
Moreover, we carefully design a more challenging test-time manipulation task which extends PnP (place), namely pick-and-where-to-place (PnWP). As shown in Fig. 4, the agent must find a global optimal trajectory given non-concave guide function $\mathcal{J}(\boldsymbol{\tau})=-\sum_{i}c_{1}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{local}}\right\rVert-c_{2}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{mid}}\right\rVert+c_{3}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{\text{global}}\right\rVert$ where $c_{1},c_{2},c_{3}$ are positive constants and $\boldsymbol{s}_{\text{local}},\boldsymbol{s}_{\text{mid}},\boldsymbol{s}_{\text{global}}$ are the local, middle, and global target positions, respectively. Both $\boldsymbol{s}_{\text{global}}$ and $\boldsymbol{s}_{\text{local}}$ are equidistant from the robot arm’s attachment point. Since the local optimum has a wide peak while the global optimum has a narrow peak, agents easily get trapped in local optima without sufficient exploration. While PnP requires fitting blocks into a target configuration, PnWP challenges planners to distinguish between globally optimal and suboptimal arrangements.
@@
We report the performance of TDP, baselines, and ablations in Table˜2. Mono-level guided sampling methods (i.e., AdaptDiffuser, MCSS(+SS), TAT, and TDP (w/o child)) tend to converge to local optima, often stacking all blocks at a single position, since a landscape of local optima is spread out in a broader range as shown in Fig. 4. In contrast, bi-level sampling approaches (i.e., TDP and TDP (w/o PG)), which combine parent branching and sub-trajectory refinement, are better able to identify globally optimal placements consistently. Notably, TDP outperforms AdaptDiffuser both on the standard benchmark (PnP) and on the custom task (PnWP). This demonstrates that TDP’s test-time scalability and generalization enable zero-shot planning that surpasses the training-per-task approach.
@@
The key advantage of our method over the baselines comes from employing unconditional PG in the parent branching phase, enabling broad exploration of the trajectory space. Both TDP and its ablations produce diverse trajectories (see Fig. 5-a); however, TDP (w/o child) often fails to find the global optimal placement due to the use of conditional PG in parent branching, which can bias samples toward local optima. The mean pairwise trajectory distance decreases when PG is combined with sub-tree expansion, as the generated sub-trajectories share segments with their corresponding parent trajectories. Unconditional PG enables the generation of diverse, gradient-free parent trajectories. When combined with sub-tree expansion guided by the objective, this bi-level strategy supports compositional solutions that successfully stack blocks into global optimal placement (see Fig. 5-b). More detailed information on the evaluation metrics among these tasks is available in Appendix˜F.
@@ 5.4 AntMaze Multi-goal Exploration
We finally evaluate test-time multi-goal exploration capability on AntMaze [fu2021d4rldatasetsdeepdatadriven], which is more challenging than Maze2D due to its high-dimensional observation space for controlling the embodied agent. Its complexity causes the pretrained diffusion planner to predict shorter horizons than the full trajectory length [dong2024cleandiffuser]. We design a new multi-goal task in which the agent must visit all goal positions in the correct order of priority, as specified by the guide function $\mathcal{J}(\boldsymbol{\tau})=\sum_{g\in\mathcal{G}}h_{g}\cdot\exp{(-\sum_{i\leq T_{\text{pred}}}\left\lVert\boldsymbol{s}_{i}-\boldsymbol{s}_{g}\right\rVert^{2}/\sigma_{g}^{2})}$ where $\mathcal{G}$ is the set of goal positions and $h_{g}$, $\sigma_{g}$ are parameters for gaussian guide function at each goal $g\in\mathcal{G}$. For example, the agent in Fig. 6 first visits goal $g_{2}$ at $t=t_{3}$. If it subsequently visits $g_{1}$, $g_{4}$, and $g_{3}$ after $t=t_{3}$, it successfully reaches all four goals in the sequence $g_{2}\rightarrow g_{1}\rightarrow g_{4}\rightarrow g_{3}$. However, two precedence rules are violated ($g_{2}\rightarrow g_{1}$, $g_{4}\rightarrow g_{3}$) while the remaining four ($g_{2}\rightarrow g_{4}$, $g_{2}\rightarrow g_{3}$, $g_{1}\rightarrow g_{4}$, and $g_{1}\rightarrow g_{3}$) are satisfied. In this scenario, the agent achieves a goal completion score of 4/4 but only 4/6 priority sequence match accuracy. Maximum accuracy of 6/6 can only be achieved by visiting all goals in the correct prioritized order.
@@ 6 Conclusion
In summary, we propose TDP, a flexible test-time planning framework that leverages a pretrained diffusion planner via a bi-level trajectory-sampling process without training. By balancing trajectory diversity and gradient-guided refinement via a branching structure of sampled trajectories, our method addresses key limitations of conventional test-time-guided planning. Empirical results across both structured and compositional manipulation tasks demonstrate consistent performance gains over existing baselines, particularly in scenarios that demand out-of-distribution generalization. Our experiments also highlight the robustness of the framework in handling non-convex, multi-objective guidance and non-differentiable constraint problems, where naïve gradient guided methods often fail.
@@
While TDP outperforms existing planning approaches across a suite of challenging test-time control tasks, our bi-level trajectory generation process incurs additional computational cost due to the expanded search in trajectory space and the pairwise trajectory distance calculations. We analyze the additional computational time required for the two PnP tasks and the PnWP task in Appendix˜G. Future work may explore more efficient search strategies or learned priors to reduce overhead while ensuring sufficient exploration and preserving planning performance.
Acknowledgements
This work was supported by IITP grant (RS-2021-II211343: AI Graduate School Program at Seoul National Univ. (5%) and RS-2025-25442338: AI Star Fellowship Support Program (Seoul National Univ.) (60%)) and NRF grant (No.2023R1A1C200781211 (35%)) funded by the Korea government (MSIT).
@@ Appendix C Proof of Proposition 1
where the conditional distribution $X_{t}|X_{0}$ in forward process is Gaussian, i.e., $\mathcal{N}(\alpha(t)x_{0},h(t)I)$ where $h(t)=1-\alpha^{2}(t)=1-\exp(-\sqrt{t})$ [guo2024gradientguidancediffusionmodels]. First we consider the case when $X_{0}\sim\mathcal{N}(0,I)$. The orthogonal gradient can be approximated as $\nabla\mathcal{J}_{\perp}(X)\approx-\frac{1}{\sigma_{2}^{2}}(X_{\perp}-w_{\perp})\mathcal{J}_{2}(X)$ by assumption. As $\mathcal{J}_{2}(X)\geq\mathbb{E}_{X_{0}}[\mathcal{J}_{2}(X_{0})]=:b_{0}>0$ and taking expectation, it is derived by linear ODE $d\mathbb{E}[X_{t,\perp}]=\left[\left(\gamma(t)-\frac{b_{0}}{\sigma_{2}^{2}}\right)\mathbb{E}[X_{t,\perp}]+\frac{b_{0}}{\sigma_{2}^{2}}w_{\perp}\right]dt$. Solving this ODE, we get lower bound of the expectation of displacement of the final state $\mathbb{E}[|X_{T,\perp}|]\geq\exp{\left(-\Phi(T,0)\right)}\left(\frac{b_{0}}{\sigma_{2}^{2}}\int_{0}^{T}\exp{\left(\Phi(s,0)\right)}ds\right)|w_{\perp}|$ where $\Phi(t,s)=\int_{s}^{t}\left(\gamma(u)-\frac{b_{0}}{\sigma_{2}^{2}}\right)du$. Since the coefficient of direction vector $w_{\perp}$ is a tractable positive constant, the final state is always off-subspace. The final state converges linearly as the given guide function satisfies the PL condition and the Lipschitz property [karimi2020linearconvergencegradientproximalgradient]. However, the global optimal solution lies in the subspace spanned by $A$, guided sample converges to local optimal solution which is orthogonal to the subspace.
Second, we consider the case when $X_{0}\sim q_{N_{f}}(\hat{X}^{T},\Sigma^{T})$. As the small perturbation is added to the unconditional sample in the subspace, the orthogonal gradient can be approximated as $\nabla\mathcal{J}_{\perp}(X)\approx-\frac{1}{\sigma_{1}^{2}}(I-AA^{\top})(X-Av_{1})\mathcal{J}_{1}(X)=-\frac{1}{\sigma_{1}^{2}}X_{\perp}\mathcal{J}_{1}(X)$. From Eq. 7, we get linear ODE $d\mathbb{E}[X_{t,\perp}]=\left[\left(\gamma(t)-\frac{1}{\sigma_{1}^{2}}\mathcal{J}_{1}(X)\right)\mathbb{E}[X_{t,\perp}]\right]dt$. As $\mathbb{E}[X_{0,\perp}]=0$, we have $\mathbb{E}[X_{T,\perp}]=0$, which means the final state is on-subspace. Similar to the previous case, the final state converges linearly to the global optimal solution lying on the subspace as the given guide function satisfies the PL condition and Lipschitz property [karimi2020linearconvergencegradientproximalgradient].
@@ I.2 AntMaze Multi-goal Exploration
TDP surpasses sequential approaches (i.e., MCTD [yoon2025montecarlotreediffusion], Diffusion-Forcing [chen2024diffusionforcingnexttokenprediction]) on the standard maze offline benchmarks [park2025ogbenchbenchmarkingofflinegoalconditioned].
@@
TDP has advantages over sequential approaches even in single-task scenarios, in terms of trajectory optimality. For instance, in standard maze benchmarks, the shortest-path trajectory between initial and goal states represents the globally optimal solution, while longer trajectories correspond to suboptimal (local optimal) solutions [chen2024simplehierarchicalplanningdiffusion]. As maze environments scale up, identifying globally optimal solutions requires evaluating combinatorially diverse trajectory candidates with varying structures.
Single-step exploration methods like MCTD optimize actions one step at a time (via guided vs. non-guided action selection), which tends to lead to locally greedy decisions [6145622] and results in longer trajectories than the shortest path. In contrast, TDP employs a multi-step exploration framework via bi-level search, enabling per-step action selection based on multi-step future rewards evaluated across diverse-branched trajectory candidates. This approach allows TDP to be able to assess long-term action consequences and avoid the local optimal path.
@@
Learned potentials [corso2024particle] offer slight improvements in both sample diversity and overall performance. The results below compare TDP using learned versus fixed potentials on the KUKA benchmarks.